The monthly salaries of the employees at Lyman Limited follow a normal distribution with an average salary of R 14 000 and a standard deviation of R 500. ,If a female employee claims to be in the top 10% of highest paid employees at the company, what must be her minimum monthly salary
1
Expert's answer
2021-09-21T02:51:50-0400

Z = (X - μ) / σ

Z critical value = 1.28 from Z reference table (for p = 0.90)

Given μ = 14000 and σ = 500,

(X - μ) / σ = (X - 14000) / 500 = 1.28

Solving for X, we find X = 14640

Therefore,  her minimum monthly salary is approximately R 14640.